Tax preparer question

You are a tax preparer. Review the expenses submitted by the client and assist your staff with the completion of itemized deductions using the new tax laws. For expenses that are not allowed, explain why these expenses are disallowed. AGI is $45,000.

 

 Client submitted expenses:

  1. 1. Taxpreparation fees from the prior year  $ 250

  2. 2. Personalproperty taxes  $ 500

  3. 3. Reimbursedhealth insurance premiums  $1,500

  4. 4. Charitablecontributions (cash only)  $ 200

  5. 5. Prescriptions    $ 150

  6. 6. Mortgageinterest   $3,850

  7. 7. Contactlens  $ 200

  8. 8. Stateincome taxes withheld   $1,675

  9. 9. Unreimbursedmedical expenses  $ 875

  10. 10. Services provided qualified non-profit   $1,000
Answer #1

The allowable and disallowed expenses have been listed with the use of following table with explanations as and where applicable.

S.No. Expense Allowed/Disallowed Explanation
1 Tax Preparation fees from the prior year $ 250 Allowed
2 Personal property taxes $500 Allowed
3 Reimbursed health insurance premiums Disallowed The premium is reimbursed by the employer. So, in no way this expenditure can be considered to be paid by the taxpayer.
4 Charitable contributions (Cash) Only $200 Allowed
5 Prescriptions $150 Disallowed Only medical expenses in excess of 10% of adjusted gross income is allowable as a deduction. In the given case, 10% of AGI (45,000*10% = $4,500) is more than total medical expenses (150 + 200 + 875) incurred by the taxpayer.
6 Mortgage interest $3,850 Allowed
7 Contact lens Disallowed Only medical expenses in excess of 10% of adjusted gross income is allowable as a deduction. In the given case, 10% of AGI (45,000*10% = $4,500) is more than total medical expenses (150 + 200 + 875) incurred by the taxpayer.
8 State income taxes withheld Allowed
9 Unreimbursed medical expenses Disallowed Only medical expenses in excess of 10% of adjusted gross income is allowable as a deduction. In the given case, 10% of AGI (45,000*10% = $4,500) is more than total medical expenses (150 + 200 + 875) incurred by the taxpayer.
10 Services provided qualified non-profit $1,000 Disallowed The deduction is allowed only with respect to out of pocket expenses (such as car expenses) incurred in providing the services which would otherwise not have been incurred. The cost towards the value of services or time spent by the taxpayer is not deductible.